zniyab1:

A 25 gram sample of a substance that's used for drug research has a k-value of 0.1229 Find the substance's half-life in days.

payload:

No direct answers

justjm:

Use the general formula for half-life \(A_t = A_oe^{-kt}\) \(A_t\) = the mass at half-life point, or \(\frac{A_o}{2}=\frac{25}{2}=12.5\) k=0.1229 \(A_o=25\) t, you must find \(12.5 = 25e^{-0.1229t}\) Use the inverse of e, ln (natural log) to find t. \(ln12.5 = ln(25e^{-0.1229t})\) \(ln12.5 = ln25-0.1229t\) Use your calculator to evaluate the value of t. That should give you your half life

payload:

sorry